OpenStudy (anonymous):

How i find a1 and a2 if friction is negligible?

OpenStudy (anonymous):

any one o explain me ...

OpenStudy (anonymous):

seriously what is that second pulley thing doing there?! i can't understand hwo the system will get accelerated in the first place!

OpenStudy (anonymous):

i got -5/6 for a1, can u tell me if it is correct

OpenStudy (anonymous):

?

OpenStudy (anonymous):

how u got -5/6 for a1 ?

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

i think a1=1/2 a2 but i dont know relation between m and a

OpenStudy (anonymous):

I got \[a _{1}=-\frac{ m_{1}g }{ 4m_{2} }\] and \[a_{2}=\frac{ m_{1}g }{ 2m_{2} }\]

OpenStudy (anonymous):

i used lagrangian method

OpenStudy (anonymous):

how u used lagranian method for this case ? can u explain

OpenStudy (anonymous):

 now u just have to express potential and kinetic energies in terms of x, then writr L=T-V, and use euler-lagrange equations